Divide and Simplify
2+5i/3-i

Answer by Alan3354(69443)

Divide and Simplify
2+5i/3-i
----------
Multiply NUM and DEN by the conjugate of the DEN.
=(3+i)*(2+5i)/(3+i)*(3-i)
= (6 + 17i - 5)/(9+1)
= (1+17i)/10